OpenStudy (anonymous):

Sooo we have y over 4 minus y over 2, and I'm lost

OpenStudy (tkhunny):

Can you subtract the fractions \(\dfrac{1}{4}-\dfrac{1}{2}\)?

OpenStudy (anonymous):

Remember you need to have a LCD on the bottom! Can you tell me what the least common denominator is?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

So really it's just 1/4y - 1/4y, the answer being No solution?

OpenStudy (anonymous):

Or, no, 1/4y - 2/4y

OpenStudy (anonymous):

Being -1/4y

OpenStudy (anonymous):

I would put that yes. But where does the Y come from?

OpenStudy (anonymous):

The question is y/4 -y/2 = -8

OpenStudy (tkhunny):

y/4 - y/2 = -8 -y/4 = -8 y/4 = 8 y = 32 Not as complicated as maybe you were thinking?

OpenStudy (anonymous):

Not at all, thank you tkhunny and ohlook...

OpenStudy (triciaal):